OpenStudy (anonymous):

A quantity of insulin used to regulate sugar in the bloodstream breaks down by about 5% each minute. A body-weight adjusted dose is generally 10 units. (a) Write a function representing the amount of the dose that remains. (b) Graph the function. (c) About how much insulin remains after 10 minutes? How do I solve this??

OpenStudy (anonymous):

Let A(t) represent the amount of dose that remains after t minutes. a) A(t) = 10 (0.95)^t c) 10 (0.95)^10 = 5.987 units d) 0.5 = 10 (0.95)^t solve for t 0.05 = (0.95)^t log (0.05) = t log(0.95) t = log(0.05) / log(0.95) = 58.4 minutes

OpenStudy (anonymous):

Thank youuuuuu!! :)